the enthalpy and entropy change for the reaction are -3.9KJ/ mol and + 56.6J/ mol K respectively at 25 degree celcius. what is the free energy change in kj/ mole? is this reaction always spontaneous never spontaneous or does it depend on the Temperature?


1
Expert's answer
2021-06-14T13:37:27-0400

∆G = ∆H - ∆TS

∆H = -3.9KJ/mol

T = 25 + 273 = 298K

∆S = 56.6J/mol = 0.056KJ/mol

∆G = -3.9 - 298(0.056)

= -3.9 - 16.688

= -20.588KJ/mol

The reaction is spontaneous